Circular No. 7684

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dr IN NGC 4932 A. Dimai, Cortina d’Ampezzo, Italy, reports the discovery of an ap-

parent supernova (mag ∼ 16.5) by A. Gabrjelcic, Agordo, Italy, on images (limiting mag ∼ 18.0) taken on Aug. 14.88 and 15.85 UT with a 0.28-m reflector (+ HiRes 1600 CCD camera) in the course of the CROSS program. The new object was also imaged at mag 16.6 with a 0.51-m reflector at Col Drusci on Aug. 16.83 (limiting mag ∼ 19.0). SN 2001dr is located at α = 13^{h}02^{m}39^{s}.10, δ = +50^{o}26′39′′.5 (equinox 2000.0), which is 9′′ east and 12′′ north of the nucleus of NGC 4932. The new object is not present on Palomar Sky Survey images.

SUPERNOVA 2001ds IN UGC 1654 The Central Bureau has received reports that independent discoveries

have been made of a new supernova in UGC 1654 by T. Boles, Coddenham, England (via G. M. Hurst, Basingstoke, England), on unfiltered CCD images taken for the U.K. Nova/Supernova Patrol on Aug. 14.999 (mag 16.2) and 16 UT (mag 16.6), and by LOTOSS (cf. IAUC 7514; via D. J. Hutchings and W. D. Li, University of California at Berkeley) on unfiltered CCD images taken with the Katzman Automatic Imaging Telescope (KAIT) on Aug. 16.5 and 17.4 (mag ∼ 18.3). M. Armstrong, Rolvenden, England, also obtained a confirming image on Aug. 16.972. The LOTOSS position for SN 2001ds is α = 2^{h}10^{m}08^{s}.73, δ = +36^{o}42′20′′.3 (equinox 2000.0), which is 4′′.1 west and 3′′.6 north of the nucleus of UGC 1654. A KAIT image taken on Jan. 17.2 showed nothing at this position (limiting mag ∼ 19.0). Boles’ position end figures, averaged from frames on two nights, are 08^{s}.80, 20′′.1, or 5′′.0 west and 3′′.5 north of the galaxy’s nucleus. SN 2001ds is absent from an image taken by Boles on 1999 Dec. 4.941 (limiting mag 18.0) and from Palomar Sky Survey images taken in 1954, 1989, and 1992.

NOVA IN M31 P. Martin and W. D. Li, University of California at Berkeley, report the

discovery of an apparent nova in M31 (= NGC 224) on unfiltered images taken by the KAIT on Aug. 14.5 (mag ∼ 17.6) and 16.5 UT (mag ∼ 16.8). The nova is located at α = 0^{h}42^{m}18^{s}.47, δ = +41^{o}12′38′′.9 (equinox 2000.0), which is 291′′ west and 210′′ south of the nucleus of M31. A KAIT image of the same field taken on Aug. 13.5 showed nothing at this position (limiting mag ∼ 19.0).

2001 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green